Earth's distance from the Sun varies from 1.47*10^
Answer #1

R1=1.47*1011

R2=1.52*1011

Mass of the Earth=5.972*1024=ME

Mass of the Sun=1.989*1030=MS

Change in potential energy=(GMSME/R1)-(GMSME/R2)

[Put value of G,MS,ME,R1&R2]

  Change in potential energy=1.765*1012J